Consider how best to prepare one liter of a buffer solution with pH = 10.93 using one of the weak acid/conjugate base systems shown here.

Weak Acid Conjugate Base Ka pKa

HC2O4-

C2O42-

6.4 x 10-5

4.19

H2PO4-

HPO42-

6.2 x 10-8

7.21

HCO3-

CO32-

4.8 x 10-11

10.32



How many grams of the sodium salt of the weak acid must be combined with how many grams of the sodium salt of its conjugate base, to produce1.00 L of a buffer that is 1.00 M in the weak base?

grams sodium salt of weak acid =

grams sodium salt of conjugate base =
Answer #1

Sol . As pKa value of HCO3- / CO32- system is 10.32 which is near to pH 10.93 . So , this weak acid / conjugate base system is preferred to prepare 1 L of buffer with pH 10.93  

Now , Molarity of weak acid = [HCO3-] = 1.00 M  

So, Using Henderson - Hasselbalch equation ,  

pH = pKa + log( [CO32-] / [HCO3-] )  

10.93 = 10.32 + log ( [CO32-] / 1.00 )  

[CO32-] / 1.00 = 100.61 = 4.0738

So , [CO32-] = 4.0738 M

Now , Molarity of sodium salt of HCO3- = [NaHCO3] = 1.00 M

and , Molarity of sodium salt of CO32- = [Na2CO3] = 4.0738 M

As Volume of buffer = 1 L

So , moles of NaHCO3 = 1.00 × 1 = 1 mol  

moles of Na2CO3 = 4.0738 × 1 = 4.0738 mol

As Molar mass of NaHCO3 = 84.007 g/mol

So , Mass of NaHCO3 = Moles of NaHCO3 × Molar mass of NaHCO3 = 1 × 84.007 = 84.007 g  

Also , Molar mass of Na2CO3 = 105.989 g/mol  

So , Mass of Na2CO3 = Moles of Na2CO3 × Molar mass of Na2CO3 = 4.0738 × 105.989 = 431.778 g